INVESTIGATION OF THE 19.5 CM^{−1} BAND IN BENZENE-D_{2}O

\underline{BRIAN S. MEEHAN}, GEOFFREY A. BLAKE, California Institute of Technology, M/S 139-74, Pasadena, CA 91125.

The spectrum of benzene-D_{2}O has been Previously Investigated At Low Resolution Using Mass-Selective, Ionization-Loss Stimulated Raman Spectroscopy Tunable Far Infrared Laser (TuFIR) Spectroscopy^{B}. Some Of The Raman Active Bands Have Been Assigned By Direct comparison to high resolution spectra. However, the nature of the band centered about 19.5 cm^{−1} remains unknown. By fitting high resolution rotational lines in this region to a Hamiltonian using l -type doubling terms, it is seen that this motion is a bending mode, perpendicular to the symmetry axis, that is at least doubly degenerate. By direct comparison of the TuFIR and Raman data, it is possible to definitively confirm this assignment.

a P. Maxton, M. Schaeffer, and P. Felker, Chem.  Phys.  Lett.  \underline{176}, 603, 1995a.

b Sakae Suzuki, Ph.D Thesis, California Institute of Technology.